state the discriminant for 2y^2 +8y =-3

Answer by AnlytcPhil(1116)

State the discriminant for
2y² + 8y = -3 

b^2-4ac   8^2-4(1) (-3)
             -360

That's altogether wrong. 

You must first get a 0 on one side and the trinomial 
in descending order on the other:

                          2y² + 8y = -3

To get zero on the right add 3 to both sides

                      2y² + 8y + 3 = 0

Compare that to       Ax² + Bx + C = 0

A=2 B=8 and C=3

Discriminant = B²-4AC = 8² - 4(2)(3) = 64 - 24 = 40

Edwin
AnlytcPhil@aol.com